OpenStudy (anonymous):

2/3x= 26 solve for variable 39 -39 17.333.. -17.333...

OpenStudy (anonymous):

multiply both sides of the equation by the reciprocal of x's coefficient. This leave x on the right side of the equation by itself.

OpenStudy (anonymous):

(3/2)(2/3)x=26(3/2) x=39